OpenStudy (anonymous):

Manny and Emma are both competing in a long jump competition at their school. Manny can jump four fifths of the distance that Emma can jump. Manny can jump 40 inches. How far can Emma jump?

OpenStudy (diyadiya):

Lets assume that Emma can jump \(x\) inches Its given in the question that Manny can jump our fifths of the distance that Emma can jump. Therefore, Manny can jump \( \Large\frac{4}{5}x \)inches But Manny can jump 40inches Hence,\[\frac{4}{5}x= 40\] @KeemoD143 Can you solve 4 \(\large x\) now?

OpenStudy (diyadiya):

*solve for

OpenStudy (anonymous):

ahhhh HMMM no :)

OpenStudy (anonymous):

is the answer 40.8??

OpenStudy (diyadiya):

\[\frac{4}{5}x = 40\]To find \(x\) we multiply both sides by 5 \[5 \times \frac{4}{5} x = 40 \times 5 \]\[\cancel{5} \times \frac{4}{\cancel{5}} x = 40 \times 5 \]\[4x= 200\] Now divide both sides by 4 you'll get the value of \(x\) that's the distance Emma can jump

OpenStudy (diyadiya):

What did you get??